OpenStudy (anonymous):

a basketball team player has made 8 free throws out of 20. if she continues at the same rate, how many shots will she make if 50 attempted?

OpenStudy (anonymous):

that person made 8/20

OpenStudy (anonymous):

8/20= B/50

OpenStudy (anonymous):

where B is the free throws when he tries 50 times

OpenStudy (anonymous):

Just solve for B

OpenStudy (anonymous):

\[\frac{8}{20} of the throw will be goaled\] so 20 will be made out of 50 throw